# Metric3D v2: A Versatile Monocular Geometric Foundation Model for Zero-shot Metric Depth and Surface Normal Estimation

### v1

@@ 3.2 Canonical Camera Transformation
The core idea is to set up a canonical camera space ($(f_{x}^{c},f_{y}^{c})$, $f_{x}^{c}=f_{y}^{c}=f^{c}$ in experiments) and transform all training data to this space. Consequently, all data can roughly be regarded as captured by the canonical camera. We propose two transformation methods, i.e. either transforming the input image ($\mathbf{I}\in\mathbb{R}^{H\times W\times 3}$) or the ground-truth (GT) label ($\mathbf{D}\in\mathbb{R}^{H\times W}$). The original intrinsics are $\{f,u_{0},v_{0}\}$.
@@
Method1: transforming depth labels (CSTM_label). Fig. 5’s ambiguity is for depths. Thus our first method directly transforms the ground-truth depth labels to solve this problem. Specifically, we scale the ground-truth depth ($\mathbf{D}^{*}$) with the ratio $\omega_{d}=\frac{f^{c}}{f}$ in training, i.e., $\mathbf{D}^{*}_{c}=\omega_{d}\mathbf{D}^{*}$. The original camera model is transformed to $\{f^{c},u_{0},v_{0}\}$. In inference, the predicted depth ($\mathbf{D}_{c}$) is in the canonical space and needs to perform a de-canonical transformation to recover the metric information, i.e., $\mathbf{D}=\frac{1}{\omega_{d}}\mathbf{D}_{c}$. Note the input $\mathbf{I}$ does not perform any transformation, i.e., $\mathbf{I}_{c}=\mathbf{I}$.
@@
Method2: transforming input images (CSTM_image). From another view, the ambiguity is caused by the similar image appearance. Thus this method is to transform the input image to simulate the canonical camera imaging effect. Specifically, the image $\mathbf{I}$ is resized with the ratio $\omega_{r}=\frac{f^{c}}{f}$, i.e., $\mathbf{I}_{c}=\mathcal{T}(\mathbf{I},\omega_{r})$, where $\mathcal{T}(\cdot)$ denotes image resize. The optical center is resized, thus the canonical camera model is $\{f^{c},\omega_{r}u_{0},\omega_{r}v_{0}\}$. The ground-truth labels are resized without any scaling, i.e., $\mathbf{D}^{*}_{c}=\mathcal{T}(\mathbf{D}^{*},\omega_{r})$. In inference, the de-canonical transformation is to resize the prediction to the original size without scaling, i.e., $\mathbf{D}=\mathcal{T}(\mathbf{D}_{c},\frac{1}{\omega_{r}})$.
@@
Fig. 9 shows the pipeline. After performing either transformation, we randomly crop a patch for training. The cropping only adjusts the FOV and the optical center, thus not causing any metric ambiguity issues. In the labels transformation method $\omega_{r}=1$ and $\omega_{d}=\frac{f^{c}}{f}$, while $\omega_{d}=1$ and $\omega_{r}=\frac{f^{c}}{f}$ in the images transformation method. During training, the transformed ground-truth depth labels $\mathbf{D}^{*}_{c}$ are then used as supervision. Notably, as surface normals do not suffer from any metric ambiguity, we impose no transformation to normal labels ${\mathbf{N}}^{*}$.
Mix-data training is an effective way to boost generalization. We collect 18 datasets for training, see Tab. V for details. In the mixed data, over 10K different cameras are included. All collected training data have included paired camera intrinsic parameters, which are used in our canonical transformation module.
@@ 3.3 Jointly optimizing depth and normal
We propose to optimize metric depth and surface normal jointly in an end-to-end manner. This optimization is primarily aimed at leveraging the large amount of annotation knowledge available in depth datasets to improve normal estimation, particularly in outdoor scenarios where depth datasets contain significantly more annotations than normal datasets. In our experiments, we collect from the community $9488$K images with depth annotations across $14$ outdoor datasets while less than $20$K outdoor normal-labeled images, presented in Tab. V.
Figure 10: Joint depth and normal optimization. In the canonical space, we deploy recurrent blocks composed of ConvGRU sub-blocks (C-RGU) and projection heads (Proj.) to predict the updates $\Delta$. During optimization, intermediate low-reolsution depth and normal $\hat{\mathbf{D}}_{c}^{0}$ $\hat{\mathbf{N}}_{u}^{0}$ are initially given by the decoder, and then iteratively refined by the predicted updates $\Delta$. After $T+1$ iterations, the optimized intermediate predictions $\hat{\mathbf{D}}_{c}^{T+1}$ $\hat{\mathbf{N}}_{u}^{T+1}$ are upsampled and post-processed to obtain the final depth $\mathbf{D}_{c}$ in the canonical space and the final normal $\mathbf{N}$.
To facilitate knowledge flow across depth and normal, we implement the learning-based optimization with recurrent refinement blocks, as depicted in Fig 10. Unlike previous monocular methods [42, 44], our method updates both depth and normal iteratively through these blocks. Inspired by RAFT [46, 45], we iteratively optimize the intermediate low-resolution depth $\hat{\mathbf{D}_{c}}$ and unnormalized normal $\hat{\mathbf{N}_{u}}$, where ^ denotes low resolution prediction $\hat{\mathbf{D}}_{c}\in\mathbb{R}^{\frac{H}{4}\times\frac{W}{4}}$, and $\hat{\mathbf{N}}_{u}\in\mathbb{R}^{\frac{H}{4}\times\frac{W}{4}\times 3}$, and the subscript c means the depth $\hat{\mathbf{D}}_{c}$ is in canonical space. As sketched in Fig. 10, $\hat{\mathbf{D}}_{c}^{t}$ and $\hat{\mathbf{N}}_{u}^{t}$ represent the low-resolution depth and normal optimized after step $t$, where $t=0,1,2,\dots,T$ denotes the step index. Initially, at step $t=0$, $\hat{\mathbf{D}}_{c}^{0}$ and $\hat{\mathbf{N}}_{u}^{0}$ are given by the decoder. In addition to updating depth and normal, the optimization module also updates hidden feature maps $\mathbf{H}^{t}$, which are initialized by the decoder. During each iteration, the learned recurrent block $\mathcal{F}$ output updates $\Delta\hat{\mathbf{D}}_{c}$, $\Delta\hat{\mathbf{N}}_{u}$ and renews the hidden features $\mathbf{H}$:
@@ 4 Experiments
Dataset details. We have curated a comprehensive dataset comprising $18$ public RGB-D datasets, totaling over $16$ million data points for training purposes. It spreads over diverse indoor and outdoor scenes. Notably, approximately $10$ million frames are annotated with normal, but the majority of these annotations pertain to indoor scenes exclusively. It’s worth mentioning that all datasets have provided camera intrinsic parameters. Apart from the test split of training datasets, we collect $7$ unseen datasets for robustness and generalization evaluation. Details of employed training and testing data are reported in Tab. V.
@@
Ours ViT-L CSTM_label ZS $\bm{0.697}$ $\bm{0.762}$ $\bm{0.788}$ $\underline{19.6}$ $\bm{5.7}$ $\underline{35.2}$
@@
Evaluation details for monocular depth and normal estimation. a) To show the robustness of our metric depth estimation method, we test on 8 zero-shot benchmarks, including NYUv2 [58], KITTI [59], NuScenes [103], iBIMS-1 [104], DIODE [105] (indoor and outdoor, and full), ETH3D [106]. Following previous works [17], absolute relative error (AbsRel), the accuracy under threshold ($\delta_{i}<1.25^{i},i=1,2,3$), root mean squared error (RMS), root mean squared error in log space (RMS_log), and log10 error (log10) metrics are employed. For KITTI and NYU benchmarks, we report the result of zero-shot and fine-tuning testing results. b) For normal estimation tasks and ablations, employ several error metrics to assess performance. Specifically, we calculate the mean (mean), median (median), and rooted mean square (RMS_normal) of the angular error as well as the accuracy under threshold of {$11.25^{\circ}$, $22.5^{\circ}$, $30.0^{\circ}$} consistent with methodologies established in previous studies  [33]. We conduct in-domain evaluation using the Scannet dataset, while the NYU and iBIMS-1 datasets are reserved for zero-shot generalization testing. c) Furthermore, we also follow current affine-invariant depth benchmarks [25, 29] (Tab. IV) to evaluate the generalization ability on $5$ zero-shot datasets, i.e., NYUv2, DIODE, ETH3D, ScanNet [107], and KITTI. We mainly compare with large-scale data trained models. Note that in this benchmark we follow existing methods to apply the scale shift alignment before evaluation.
@@
Evaluation details for reconstruction and SLAM. a) To evaluate our metric 3D reconstruction quality, we randomly sample 9 unseen scenes from NYUv2 and use colmap [108] to obtain the camera poses for multi-frame reconstruction. Chamfer $l_{1}$ distance and the F-score [109] are used to evaluate the reconstruction accuracy. b) In dense-SLAM experiments, following Li et al. [110], we test on the KITTI odometry benchmark [59] and evaluate the average translational RMS drift (%$\%,t_{rel}$) and rotational RMS drift (${}^{\circ}/100m,r_{rel}$) errors [59].
@@
Generalization over diverse scenes. Affine-invariant depth benchmarks decouple the scale’s effect, which aims to evaluate the model’s generalization ability to diverse scenes. Recent impact works, such as MiDaS, LeReS, and DPT, achieved promising performance on them. Following them, we test on 5 datasets and manually align the scale and shift to the ground-truth depth before evaluation. Results are reported in Tab. IV. Although our method enforces the network to recover more challenging metric information, our method outperforms them by a large margin on all datasets.
@@
Evaluation on surface normal benchmarks. We evaluate our methods on ScanNet, NYU, and iBims-1 surface normal benchmarks. Results are reported in Tab. II. Firstly, we organize a zero-shot testing benchmark on NYU dataset, see methods denoted with ‘ZS’ in the table. We compare with existing methods which are trained on ScanNet or Taskonomy and have achieved promising performance on them, such as Polymax [34] and Bae et al. [33]. Our method surpasses them over most metrics. Then, comparing with methods that have been overfitted the NYU data domain for hundreds of epochs (see methods marked with blue), our zero-shot testing outperforms them on all metrics. Our post-finetuned models (see ‘FT’ marks) further boost the performance. Similarly, we also achieve SoTA performance on iBims-1 and Scannet benchmarks. For the iBims-1 dataset, we follow IronDepth[42] to generate the ground-truth normal annotations.
@@ 4.1 Zero-shot Generalization
TABLE VI: Quantitative comparison of 3D scene reconstruction with LeReS [25], DPT [28], RCVD [128], SC-DepthV2 [129], and a learning-based MVS method (DPSNet [48]) on 9 unseen NYUv2 scenes. Apart from DPSNet and ours, other methods have to align the scale with ground truth depth for each frame. As a result, our reconstructed 3D scenes achieve the best performance.
@@ 4.2 Applications Based on Our Method
In these experiments, we apply the CSTM_image model to various tasks.
Figure 13: Reconstruction of zero-shot scenes with multiple views. We sample several NYUv2 scenes for 3D reconstruction comparison. As our method can predict accurate metric depth, thus all frame’s predictions are fused together for scene reconstruction. By contrast, LeReS [25]’s depth is up to an unknown scale and shift, which causes noticeable distortions. DPSNet [48] is a multi-view stereo method, which cannot work well on low-texture regions.
3D scene reconstruction . To demonstrate our work can recover the 3D metric shape in the wild, we first do the quantitative comparison on 9 NYUv2 scenes, which are unseen during training. We predict the per-frame metric depth and then fuse them together with provided camera poses. Results are reported in Tab. VI. We compare with the video consistent depth prediction method (RCVD [128]), the unsupervised video depth estimation method (SC-DepthV2 [129]), the 3D scene shape recovery method (LeReS [25]), affine-invariant depth estimation method (DPT [28]), and the multi-view stereo reconstruction method (DPSNet [48]). Apart from DPSNet and our method, other methods have to align the scale with the ground truth depth for each frame. Although our method does not aim for the video or multi-view reconstruction problem, our method can achieve promising consistency between frames and reconstruct much more accurate 3D scenes than others on these zero-shot scenes. From the qualitative comparison in Fig. 13. our reconstructions have much less noise and outliers.
Dense-SLAM mapping. Monocular SLAM is an important robotics application. It only relies on a monocular video input to create the trajectory and dense 3D mapping. Owing to limited photometric and geometric constraints, existing methods face serious scale drift problems in large scenes and cannot recover the metric information. Our robust metric depth estimation method is a strong depth prior to the SLAM system. To demonstrate this benefit, we naively input our metric depth to the SoTA SLAM system, Droid-SLAM [37], and evaluate the trajectory on KITTI. We do not do any tuning on the original system. Trajectory comparisons are reported in Tab. VII. As Droid-SLAM can access accurate per-frame metric depth, like an RGB-D SLAM, the translation drift ($t_{rel}$) decreases significantly. Furthermore, with our depths, Droid-SLAM can perform denser and more accurate 3D mapping. An example is shown in Fig. 3 and more cases are shown in the supplementary materials.
We also test on the ETH3D SLAM benchmarks. Results are reported in Tab. VIII. Droid with our depths has much better SLAM performance. As the ETH3D scenes are all small-scale indoor scenes, the performance improvement is less than that on KITTI.
@@
Metrology in the wild. To show the robustness and accuracy of our recovered metric 3D, we download Flickr photos captured by various cameras and collect coarse camera intrinsic parameters from their metadata. We use our CSTM_image model to reconstruct their metric shape and measure structures’ sizes (marked in red in Fig. 14), while the ground-truth sizes are in blue. It shows that our measured sizes are very close to the ground-truth sizes.

### v4

@@ 3.2 Canonical Camera Transformation
The fundamental concept entails establishing a canonical camera space ($(f_{x}^{c},f_{y}^{c})$, with $f_{x}^{c}=f_{y}^{c}=f^{c}$ in experimental settings) and transposing all training data into this designated space. Consequently, all data can be broadly construed as being captured by the canonical camera. We propose two transformation methods, i.e. either transforming the input image ($\mathbf{I}\in\mathbb{R}^{H\times W\times 3}$) or the ground-truth (GT) label ($\mathbf{D}\in\mathbb{R}^{H\times W}$). The initial intrinsics are $\{f,u_{0},v_{0}\}$.
@@
Method1: transforming depth labels (CSTM_label). Fig. 5’s ambiguity is for depths. Consequently, our initial approach directly addresses this issue by transforming the ground-truth depth labels. Specifically, we rescale the ground-truth depth ($\mathbf{D}^{*}$) using the ratio $\omega_{d}=\frac{f^{c}}{f}$ during training, denoted as $\mathbf{D}_{c}=\omega_{d}\mathbf{D}^{*}$. The original camera model undergoes transformation to ${f^{c},u_{0},v_{0}}$. In inference, the predicted depth ($\mathbf{D}_{c}$) exists in the canonical space and necessitates a de-canonical transformation to restore metric information, expressed as $\mathbf{D}=\frac{1}{\omega_{d}}\mathbf{D}_{c}$. It is noteworthy that the input $\mathbf{I}$ remains unaltered, represented as $\mathbf{I}_{c}=\mathbf{I}$.
@@
Method2: transforming input images (CSTM_image). From an alternate perspective, the ambiguity arises due to the resemblance in image appearance. Consequently, this methodology aims to alter the input image to emulate the imaging effects of the canonical camera. Specifically, the image $\mathbf{I}$ undergoes resizing using the ratio $\omega_{r}=\frac{f^{c}}{f}$, denoted as $\mathbf{I}{c}=\mathcal{T}(\mathbf{I},\omega_{r})$, where $\mathcal{T}(\cdot)$ signifies image resizing. As a result of resizing the optical center, the canonical camera model becomes ${f^{c},\omega_{r}u_{0},\omega_{r}v_{0}}$. The ground-truth labels are resized without scaling, represented as $\mathbf{D}^{*}_{c}=\mathcal{T}(\mathbf{D}^{*},\omega_{r})$. In inference, the de-canonical transformation involves resizing the prediction to its original dimensions without scaling, expressed as $\mathbf{D}=\mathcal{T}(\mathbf{D}_{c},\frac{1}{\omega_{r}})$.
While similar transformations have been employed in MPSD [89] to normalized depth prediction, our approaches apply these modules to predict metric depth directly.
@@
Figure 9 shows the pipeline. After adopting either transformation, a patch is randomly cropped for training purposes. This cropping operation solely adjusts the field of view (FOV) and the optical center, thus averting any potential metric ambiguity issues. In the labels transformation approach, $\omega_{r}=1$ and $\omega_{d}=\frac{f^{c}}{f}$, while in the images transformation method, $\omega_{d}=1$ and $\omega_{r}=\frac{f^{c}}{f}$. Throughout the training process, the transformed ground-truth depth labels $\mathbf{D}^{*}_{c}$ are employed as supervision. Importantly, since surface normals are not susceptible to metric ambiguity, no transformation is applied to normal labels ${\mathbf{N}}^{*}$.
@@ 3.3 Jointly optimizing depth and normal
We propose to optimize metric depth and surface normal jointly in an end-to-end manner. This optimization is primarily aimed at leveraging a large amount of annotation knowledge available in depth datasets to improve normal estimation, particularly in outdoor scenarios where depth datasets contain significantly more annotations than normal datasets. In our experiments, we collect from the community $9488$K images with depth annotations across $14$ outdoor datasets while less than $20$K outdoor normal-labeled images, presented in Tab. V.
Figure 10: Joint depth and normal optimization. In the canonical space, we deploy recurrent blocks composed of ConvGRU sub-blocks (C-RGU) and projection heads (Proj.) to predict the updates $\Delta$. During optimization, intermediate low-resolution depth and normal $\hat{\mathbf{D}}_{c}^{0}$ $\hat{\mathbf{N}}_{u}^{0}$ are initially given by the decoder, and then iteratively refined by the predicted updates $\Delta$. After $T+1$ iterations, the optimized intermediate predictions $\hat{\mathbf{D}}_{c}^{T+1}$ $\hat{\mathbf{N}}_{u}^{T+1}$ are upsampled and post-processed to obtain the final depth $\mathbf{D}_{c}$ in the canonical space and the final normal $\mathbf{N}$.
To facilitate knowledge flow across the depth and normal, we implement the learning-based optimization with recurrent refinement blocks, as depicted in Fig 10. Unlike previous monocular methods [42, 44], our method updates both depth and normal iteratively through these blocks. Inspired by RAFT [46, 45], we iteratively optimize the intermediate low-resolution depth $\hat{\mathbf{D}_{c}}$ and unnormalized normal $\hat{\mathbf{N}_{u}}$, where ^ denotes low resolution prediction $\hat{\mathbf{D}}_{c}\in\mathbb{R}^{\frac{H}{4}\times\frac{W}{4}}$, and $\hat{\mathbf{N}}_{u}\in\mathbb{R}^{\frac{H}{4}\times\frac{W}{4}\times 3}$, and the subscript c means the depth $\hat{\mathbf{D}}_{c}$ is in canonical space. As sketched in Fig. 10, $\hat{\mathbf{D}}_{c}^{t}$ and $\hat{\mathbf{N}}_{u}^{t}$ represent the low-resolution depth and normal optimized after step $t$, where $t=0,1,2,\dots,T$ denotes the step index. Initially, at step $t=0$, $\hat{\mathbf{D}}_{c}^{0}$ and $\hat{\mathbf{N}}_{u}^{0}$ are given by the decoder. In addition to updating depth and normal, the optimization module also updates hidden feature maps $\mathbf{H}^{t}$, which are initialized by the decoder. During each iteration, the learned recurrent block $\mathcal{F}$ output updates $\Delta\hat{\mathbf{D}}_{c}$, $\Delta\hat{\mathbf{N}}_{u}$ and renews the hidden features $\mathbf{H}$:
@@ 4 Experiments
Dataset details. We have meticulously assembled a comprehensive dataset incorporating 16 publicly available RGB-D datasets, comprising a cumulative total of over 16 million data points specifically intended for training purposes. This dataset encompasses a diverse array of both indoor and outdoor scenes. Notably, approximately 10 million frames within the dataset are annotated with normals, with a predominant focus on annotations relating to indoor scenes. It is noteworthy to highlight that all datasets have provided camera intrinsic parameters. Additionally, beyond the test split of training datasets, we have procured 7 previously unobserved datasets to facilitate robustness and generalization evaluations. Detailed descriptions of the utilized training and testing data are provided in Table V.
@@
Ours ViT-g CSTM_label ZS $\bm{0.697}$ $\bm{0.762}$ $\bm{0.788}$ $\underline{19.6}$ $\bm{5.7}$ $35.2$
@@
Evaluation details for monocular depth and normal estimation. a) To demonstrate the robustness of our metric depth estimation method, we evaluate on 7 zero-shot benchmarks, including NYUv2, KITTI [106], ScanNet [107], NuScenes [108], iBIMS-1[109], and DIODE [110] (both indoor and outdoor). Following previous studies, we use metrics such as absolute relative error (AbsRel), accuracy under threshold ($\delta_{i}<1.25^{i},i=1,2,3$), root mean squared error (RMS), root mean squared error in log space (RMS_log), and log10 error (log10). We report results for zero-shot and fine-tuning testing on the KITTI and NYU benchmarks. b) For normal estimation tasks and ablations, employ several error metrics to assess performance. Specifically, we calculate the mean (mean), median (median), and rooted mean square (RMS_normal) of the angular error as well as the accuracy under threshold of {$11.25^{\circ}$, $22.5^{\circ}$, $30.0^{\circ}$} consistent with methodologies established in previous studies  [33]. We conduct in-domain evaluation using the Scannet dataset, while the NYU and iBIMS-1 datasets are reserved for zero-shot generalization testing. c) Furthermore, we also follow current affine-invariant depth benchmarks [25, 29] (Tab. IV) to evaluate the generalization ability on $5$ zero-shot datasets, i.e., NYUv2, DIODE, ETH3D, ScanNet [107], and KITTI. We mainly compare with large-scale data trained models. Note that in this benchmark we follow existing methods to apply the scale shift alignment before evaluation.
@@
[…]
@@
Generalization over diverse scenes. Affine-invariant depth benchmarks decouple the scale’s effect, which aims to evaluate the model’s generalization ability to diverse scenes. Recent impact works, such as MiDaS, LeReS, DPT, Marigold, and DepthAnything achieved promising performance on them. Following them, we test on 5 datasets and manually align the scale and shift to the ground-truth depth before evaluation. Results are reported in Tab. IV. Although our method enforces the network to recover the more challenging metric, our method outperforms them on all datasets.
@@
Evaluation on surface normal benchmarks. We evaluate our methods on ScanNet, NYU, and iBims-1 surface normal benchmarks. Results are reported in Tab. II. Firstly, we organize a zero-shot testing benchmark on NYU dataset, see methods denoted with ‘ZS’ in the table. We compare with existing methods which are trained on ScanNet or Taskonomy and have achieved promising performance on them, such as Polymax [34] and Bae et al. [33]. Our method surpasses them over most metrics. Cmparing with methods that have been overfitted the NYU data domain for hundreds of epochs (marked with blue), our zero-shot testing outperforms them on all metrics. Our post-finetuned models (‘FT’ marks) further boost the performance. Similarly, we also achieve SoTA performance on iBims-1 and Scannet benchmarks. For the iBims-1 dataset, we follow IronDepth[42] to generate the ground-truth normal annotations.
@@ 4.1 Zero-shot Generalization
TABLE VI: Quantitative comparison of 3D scene reconstruction with LeReS [25], DPT [28], RCVD [130], SC-DepthV2 [131], and two learning-based MVS methods (DPSNet [48], SimpleRecon [132]) on 9 unseen NYUv2 scenes. Apart from the MVS approaches and ours, other methods have to align the scale with ground truth depth for each frame. As a result, our reconstructed 3D scenes achieve the best performance.
@@ 4.2 Applications Based on Our Method
We apply the CSTM_image model to various tasks.
Figure 13: Reconstruction of zero-shot scenes with multiple views. We sample several NYUv2 scenes for 3D reconstruction comparison. As our method can predict accurate metric depth, thus all frame’s predictions are fused directly for reconstruction. By contrast, LeReS [25]’s depth is up to an unknown scale and shift, causing noticeable distortions. For MVS methods, DPSNet [48] fails on low-texture backgrounds, while SimpleRecon [132] distorts regions without sufficient observations.
3D scene reconstruction . To present our method’s ability to recover real-world metric 3D, we first conduct a quantitative comparison on 9 unseen NYUv2 scenes. We predict per-frame metric depth and fuse these with the provided camera poses, with results detailed in Table VI. We compare our approach to several methods: the video consistent depth prediction method (RCVD [130]), unsupervised video depth estimation (SC-DepthV2 [131]), 3D scene shape recovery (LeReS [25]), affine-invariant depth estimation (DPT [28]), and multi-view stereo reconstruction (DPSNet [48], SimpleRecon [132]). Except for the multi-view approaches and our method, all others require aligning scales with ground truth depth for each frame. While our approach is not specifically designed for video or multi-view reconstruction, it demonstrates promising frame consistency and significantly more accurate 3D scene reconstructions in these zero-shot scenarios. Qualitative comparisons in Fig. 13 reveal that our reconstructions exhibit considerably less noise and fewer outliers.
Dense-SLAM mapping. Monocular SLAM is a key robotic application that uses a single video input to create trajectories and dense 3D maps. However, due to limited photometric and geometric constraints, existing methods struggle with scale drift in large scenes and fail to recover accurate metric information. Our robust metric depth estimation serves as a strong depth prior for the SLAM system. To demonstrate this, we input our metric depth into the state-of-the-art SLAM system, Droid-SLAM [37], and evaluate the trajectory on KITTI without any tuning. Results are shown in Table VII. With access to accurate per-frame metric depth, Droid-SLAM experiences a significant reduction in translation drift ($t_{rel}$). Additionally, our depth data enables Droid-SLAM to achieve denser and more precise 3D mapping, as illustrated in Fig. 3 and detailed in the supplementary materials.
We also tested on the ETH3D SLAM benchmarks, with results in Table VIII. Using our metric depth predictions, Droid-SLAM shows improved performance, although the gains are less pronounced in the smaller indoor scenes of ETH3D compared to KITTI.
@@
Metrology in the wild. To demonstrate the robustness and accuracy of our recovered metric 3D shapes, we downloaded Flickr photos taken by various cameras and extracted coarse camera intrinsic parameters from their metadata. We utilized our CSTM_image model to reconstruct the metric shapes and measure the sizes of structures (marked in red in Fig. 14), with ground-truth sizes shown in blue. The results indicate that our measured sizes closely align with the ground-truth values.